# First Step Towards Matter Power Spectrum Reconstruction with Stage-III Weak Gravitational Lensing Surveys

### v1

@@ Abstract
Aims. To investigate the cosmic shear measurements of Stage-III WL surveys, we compare their tomographic data by assuming a simple parametric model for the matter power spectrum. The model allows the comparison of surveys with different characteristics and, in an agnostic approach, gives insight into the shape of the matter power spectrum preferred by the data without assuming a cosmological model.
Methods. For the matter power spectrum, we assume a double power law model in scale factor and wavenumber. The best-fit amplitude and exponents are inferred in an MCMC analysis. We identify the scales to which the data is most sensitive. We test the sensitivity to different assumptions of the intrinsic alignment strength.
Results. We find that Stage-III surveys’ constraining power on the power spectrum shape and evolution is still limited. Most information can be summarised as an overall amplitude at a pivot point in wave number and scale factor, while constraints on the power law indices are considerably weaker. Nevertheless, all surveys prefer a weaker rate of growth from $z=$ 0.5 to 0.1 than predicted. The assumed intrinsic alignment strength is found to have no significant impact on the measured parameters and goodness of fit.
Conclusions. Direct estimates of the matter power spectrum from Stage-III weak lensing surveys can, in principle, be used to locate the physical origin of the observed $S_{8}$ tension. We present a simple methodology for the first steps in this direction but find that current constraints are still weak.
@@ 1 Introduction
The concordance model of cosmology, the $\Lambda$ cold dark matter ($\Lambda$CDM) model, can fit an extensive range of observations with great accuracy (see e.g. Lahav & Liddle, 2022). Examples are the fluctuations of the cosmic microwave background (CMB, Planck Collaboration et al., 2020a, Planck), baryon acoustic oscillations (BAO, DESI Collaboration et al., 2024a), and cosmic shear, the slight distortion of galaxy images by weak gravitational lensing (WL) of the large-scale structure of the Universe (Kilbinger, 2015). With the increasing accuracy of cosmological surveys, several tensions between different probes have emerged. Most notably, the tension in the value of the Hubble constant ($H_{\mathrm{0}}$) inferred from high-redshift Planck CMB measurements or local distance ladder measurements (Riess et al., 2021), and the value of $S_{8}\equiv\sigma_{\mathrm{8}}\sqrt{\Omega_{\mathrm{m}}/0.3}$. [^1] Where $\Omega_{\mathrm{m}}$ is the matter density parameter, and $\sigma_{8}$ the amplitude of the linear theory matter power spectrum parametrized as the root mean squared mass density fluctuation in spheres of radius 8 $h^{-1}$ Mpc, both at $z=0$. Here, the value of the Planck CMB measurements of the early Universe is in slight tension with the value measured from cosmic shear measurement of the local Universe (Hildebrandt et al., 2017; Asgari et al., 2021; Secco et al., 2022; Dark Energy Survey and Kilo-Degree Survey Collaboration et al., 2023; Li et al., 2023). The main proposed solutions to solving the $S_{\mathrm{8}}$ tension involve unrecognized systematics or new physics (see e.g. Lucca, 2021; Tanimura et al., 2023; Joseph et al., 2023).
@@
The tension between the probes may have different physical origins. The CMB measures linear scales and the early Universe, whereas WL probes lower redshifts and non-linear scales. Assuming a physical origin, the tension may originate either from differences in scale (high vs. low wavenumber; $k$) or in time (high vs. low redshift; $z$).
@@
Amon & Efstathiou (2022) explore the first option, introducing a one-parameter phenomenological model that modifies the power spectrum by fixing it at linear scales to the Planck prediction and predicting the non-linear suppression required to reconcile cosmic shear measurements from the Kilo-Degree Survey (KiDS) with Planck. They remain agnostic as to what causes the suppression: obvious candidates are misinterpreted baryonic components or new physics. However, the model predicts the suppression of the non-linear power spectrum using only dark matter (DM), and it does not contain any baryonic-informed components. Although baryonic suppression and non-linear collapse manifest themselves in roughly the same wavenumber regime, the suppression they model is on slightly larger scales and has a different shape than predicted by state-of-the-art cosmological hydrodynamical simulation calibrated to primarily X-ray data (see e.g. Fig. 1 of Bigwood et al., 2024).
@@
Preston et al. (2023) extended the analysis to include the most recent cosmic shear measurement from the Dark Energy Survey (DES), finding they require a somewhat lower suppression to resolve the tension with Planck. However, current measurements are not precise enough to constrain the baryonic feedback models. Recent findings suggest the simulations calibrated to X-ray data, which probe the central regions of haloes, underpredict baryonic feedback as they cannot reproduce Sunyaev-Zel’dovich (SZ) measurements, which are more sensitive to the outskirts of haloes (Amodeo et al., 2021; Bigwood et al., 2024; Hadzhiyska et al., 2024). Although it might help resolve the $S_{8}$ tension, it raises a different problem because the X-ray and SZ measurements are no longer consistent with the same baryonic feedback model.
@@
Figure 1: Observational cosmological probes of the matter power spectrum and the wavenumber ($k)$, scale factor ($a$), and redshift ($z$) ranges to which they are approximately sensitive. From early to late Universe: the cosmic microwave background (CMB, purple) CMB lensing ($\Phi\Phi$, light blue), the Lyman-alpha forest (Ly $\alpha$, orange), Quasi-Stellar Objects (QSOs, red), galaxy clustering ($gg$, green), cosmic shear ($\gamma\gamma$, dark blue), and cluster abundance (CA, white). The black curve separates the linear and non-linear collapse regimes. Constraining the matter power spectrum from different probes will allow the localization of discrepancies in the two-dimensional ($k,a$) plane and provide insight into the physical origin of tensions between the measurements of different probes.
@@
In principle, we can distinguish tensions between measurements at different scales and times by directly constraining the matter power spectrum in the wavenumber ($k$) and scale factor ($a$) plane. Fig. 1 illustrates the sensitivity of different observational cosmological probes in the ($k,a$) plane. The different probes are cosmic shear ($\gamma\gamma$), galaxy clustering ($gg$, e.g. Reid et al., 2010; Gil-Marín et al., 2016), CMB, CMB lensing ($\Phi\Phi$, e.g. Planck Collaboration et al., 2020b), cluster abundance (CA, e.g Tegmark et al., 2004), Quasi-Stellar Objects (QSOs, e.g. DESI Collaboration et al., 2024b), and the Lyman alpha forest (Ly $\alpha$, e.g. Palanque-Delabrouille et al., 2015; Karaçaylı et al., 2024). The squares correspond to probes of which their emitted radiation is directly observed, whereas the ellipses correspond to processes happening along the line of sight between emission and observation. The black curve separates the linear and non-linear collapse regimes. The estimate is obtained from class (Blas et al., 2011) by determining the wavenumber for which the non-linear estimate differs more than 5% from the linear prediction. The difference between the different cosmic times and scales probed by the CMB and cosmic shear, as described before, is clearly visible in the Figure. Comparing different probes and their cross-correlations allows for the localization of the tension, which should inform about the origin of the observed discrepancies.
@@
Preston et al. (2024) propose constraining the matter power spectrum from cosmic shear surveys. They estimate the constraining power of the matter power spectrum using synthetic cosmic shear measurements for the final data release of Stage-III and Stage-IV WL surveys. They find that while current cosmic shear surveys should only marginally be able to distinguish differences from a $\Lambda$CDM dark matter-only (DMO) model, next-generation surveys should provide clear constraints on the shape of the matter power spectrum at (mildly) non-linear scales.
@@
Here, as a first step, we infer the best-fitting shape of the matter power spectrum using the latest cosmic shear measurement of Stage-III WL surveys. We assume a simple parametric model that allows us to quantify the evolution of the matter power spectrum with $k$ and $a$ preferred by current data in an agnostic approach. The model allows the comparison of different surveys with different survey properties and does not assume a cosmological model for the matter power spectrum. We identify the scales to which the data is most sensitive and test the impact of the assumed intrinsic alignment (IA) on the goodness of fit.
@@
In Sect. 2, we summarize the relevant WL theory to estimate the measured cosmic shear two-point statistics from the matter power spectrum. Section 3 describes the data, and the assumed parametric model is introduced in Sect. 4. In Sect. 5, we present the measurements of the best-fit parametric model of the matter power spectrum. We also compare our results to the best-fit $\Lambda$CDM predictions. The main findings are summarized in Sect. 6.
@@ 2 Theory
In this section, we briefly summarize the theoretical framework for cosmic shear measurements as projected statistics of the matter power spectrum.
Cosmic shear measures the systematic distortion of galaxy images from WL by the large-scale structure. The measured signal is comprised of the signal induced by weak gravitational lensing (G), the intrinsic alignment of the galaxies (I), and the cross-correlation of the two effects. The observed cosmic shear angular power spectrum ($\mathcal{C}^{(ij)}_{\epsilon\epsilon}$) between two populations of galaxies in different tomographic bins $i$ and $j$, as a function of multipole moment $\ell$, is given by
@@
Current cosmic shear surveys use $B$-mode statistics as null tests, finding no evidence for non-zero $B$-modes or not using the range where the $B$-modes are non-zero due to unrecognized systematics. Here, we assume the $B$-modes are 0 and only use the $E$-mode signal such that $\mathcal{C}^{(ij)}_{\epsilon\epsilon}(\ell)=\mathcal{C}^{(ij)}_{\epsilon\epsilon,\mathrm{E}}(\ell)$. Assuming the modified Limber approximation, the angular power spectra are estimated from the matter power spectrum ($P_{\mathrm{m}}$) as (Limber, 1953; LoVerde & Afshordi, 2008)
@@
where $f_{\mathrm{K}}(\chi)$ is the angular diameter distance, $\chi$ the comoving distance, $z$ the redshift, and $\mathrm{a,b}\in{\mathrm{I},\mathrm{G}}$. The integral runs until the comoving line of sight horizon that is the edge of the galaxy survey ($\chi_{\mathrm{hor}}$). The kernels $W$ are given by (Kaiser, 1992)
@@
where $c$ is the speed of light, $a=1/(1+z)$ the scale factor, and $n_{\mathrm{s}}(\chi)$ is the comoving distance source distribution that relates to the source redshift distribution as $n_{\mathrm{s}}(z)\mathrm{d}z=n_{\mathrm{s}}(\chi)\mathrm{d}\chi$, and (Bridle & King, 2007)
@@
with $A_{\mathrm{IA}}$ the IA amplitude, $z_{\mathrm{piv}}$ a finite redshift pivot point, $\eta_{\mathrm{IA}}$ the redshift dependence exponent which we set to 0, $C_{1}\rho_{\mathrm{cr}}\ =0.0134$, and $D$ the linear growth factor. We assume a flat Universe such that $f_{\mathrm{K}}(\chi)=\chi$. We either adopt $A_{\mathrm{IA}}=0$ or 1 to quantify whether including intrinsic alignment better fits the data.
2.1 Cosmic shear observables
We use three different cosmic shear observables in our analysis. Each observable is a projection of the angular power spectra but is sensitive to different scales as they depend on different filter functions. The first observable is the shear two-point correlation functions (2PCFs, $\xi_{\mathrm{+/-}}$);
$\xi_{\mathrm{+/-}}^{(ij)}(\theta)=\int_{0}^{\infty}\frac{\mathrm{d}\ell\ell}{2\pi}\,\mathcal{J}_{0/4}(\ell\theta)\,\mathcal{C}^{(ij)}_{\epsilon\epsilon}(\ell),$ (5)
where $\mathcal{J}_{0}$ and $\mathcal{J}_{4}$ are cylindrical Bessel functions of the first kind of the zeroth and fourth order, respectively. Secondly, we consider Complete Orthogonal Sets of E/B-Integrals (COSEBIs). An advantage of the COSEBIs is that close to all cosmological information is contained in the first few $n$-modes, thus reducing the necessary computational time (Asgari et al., 2012). The COSEBIs ($E_{n}$) are estimated using (Schneider et al., 2010)
@@ 2.1 Cosmic shear observables
where the weight function, $W_{n}(\ell)$, is a Hankel transforms given by Equation 9 of Asgari et al. (2021). Similar to Asgari et al. (2021), we use the first 5 COSEBIs $n$-modes.
Finally, we consider band power spectra ($\mathcal{C}_{\mathrm{E},l}$, band powers). This statistic is an angular average of the two-point correlation function that has a weak correlation between values at different wavenumbers and is given by (Schneider et al., 2002)
@@
where the normalisation $\mathcal{N}_{l}$ and kernels $W_{\mathrm{EE}}$ are given by equations 18 and 26 of Joachimi et al. (2021), respectively. We assume the best-fit $\Lambda$CDM Planck cosmology for the values of the cosmological parameters and to compute the redshift-comoving distance relation and linear growth factor.

### v2

# First step toward matter power spectrum reconstruction with Stage III weak gravitational lensing surveys

@@ Abstract
Aims. To investigate the cosmic shear measurements of Stage III WL surveys, we compared their tomographic data by assuming a simple parametric model for the matter power spectrum. The model allows the comparison of surveys with different characteristics and, in an agnostic approach, gives insight into the shape of the matter power spectrum preferred by the data without assuming a cosmological model.
Methods. For the matter power spectrum, we assumed a double power-law model in scale factor and wavenumber. The best-fitting amplitude and exponents were inferred in a Markov chain Monte Carlo (MCMC) analysis. We identified the scales to which the data is most sensitive. We tested the sensitivity to different assumptions of the intrinsic alignment strength.
Results. We find that the constraining power of Stage III surveys on the power spectrum shape and evolution is still limited. Most information can be summarized as an overall amplitude at a pivot point in wavenumber and scale factor, while constraints on the power-law indices are considerably weaker. Nevertheless, all surveys show a weaker rate of growth from $z=$ 0.5 to 0.1 than predicted. The assumed intrinsic alignment strength is found to have no significant impact on the measured parameters and goodness of fit.
Conclusions. Direct estimates of the matter power spectrum from Stage III weak lensing surveys can, in principle, be used to locate the physical origin of the observed $S_{8}$ tension. We present a simple methodology for the first steps in this direction, but find that current constraints are still weak.
@@ 1 Introduction
The concordance model of cosmology, the Lambda-cold dark matter ($\Lambda$CDM) model, can fit an extensive range of observations with great accuracy (see, e.g., Lahav & Liddle, 2022). Examples are the fluctuations in the cosmic microwave background (CMB, Planck Collaboration et al., 2020a, Planck); baryon acoustic oscillations (BAOs; DESI Collaboration et al., 2024a); and cosmic shear, the slight distortion of galaxy images by weak gravitational lensing (WL) of the large-scale structure of the Universe (Kilbinger, 2015). With the increasing accuracy of cosmological surveys, several tensions between different probes have emerged. Most notably, the tension in the value of the Hubble constant ($H_{\mathrm{0}}$) inferred from high-redshift Planck CMB measurements or local distance ladder measurements (Riess et al., 2021), and the value of $S_{8}\equiv\sigma_{\mathrm{8}}\sqrt{\Omega_{\mathrm{m}}/0.3}$. [^1] Where $\Omega_{\mathrm{m}}$ is the matter density parameter and $\sigma_{8}$ the amplitude of the linear theory matter power spectrum parameterized as the root mean squared mass density fluctuation in spheres of radius 8 $h^{-1}$ Mpc, both at $z=0$. Here the value of the Planck CMB measurements of the early Universe is in slight tension with the value measured from cosmic shear measurement of the local Universe (Hildebrandt et al., 2017; Asgari et al., 2021; Secco et al., 2022; Dark Energy Survey and Kilo-Degree Survey Collaboration et al., 2023; Li et al., 2023). The main proposed solutions to solving the $S_{\mathrm{8}}$ tension involve unrecognized systematics or new physics (see, e.g., Lucca, 2021; Tanimura et al., 2023; Joseph et al., 2023).
@@
The tension between the probes may have different physical origins. The CMB measures linear scales and the early Universe, whereas WL probes lower redshifts and nonlinear scales. Assuming a physical origin, the tension may originate from differences in either scale (high vs. low wavenumber; $k$) or time (high vs. low redshift; $z$).
@@
Amon & Efstathiou (2022) explore the first option, introducing a one-parameter phenomenological model that modifies the power spectrum by fixing it at linear scales to the Planck prediction and predicting the nonlinear suppression required to reconcile cosmic shear measurements from the Kilo-Degree Survey (KiDS) with Planck. They remain agnostic as to what causes the suppression: obvious candidates are misinterpreted baryonic components or new physics. However, the model predicts the suppression of the nonlinear power spectrum using only dark matter (DM), and it does not contain any baryonic-informed components. Although baryonic suppression and nonlinear collapse manifest themselves in roughly the same wavenumber regime, the suppression they model is on slightly larger scales and has a different shape than predicted by state-of-the-art cosmological hydrodynamical simulation calibrated primarily to X-ray data (see, e.g., Fig. 1 of Bigwood et al. 2024, and, e.g., McCarthy et al., 2017; Pakmor et al., 2023; Schaye et al., 2023; Schaller et al., 2024).
@@
Preston et al. (2023) extended the analysis to include the most recent cosmic shear measurement from the Dark Energy Survey (DES), and found that they require a somewhat lower suppression to resolve the tension with Planck. However, the current measurements are not precise enough to constrain the baryonic feedback models. Recent findings suggest that the simulations calibrated to X-ray data, which probe the central regions of haloes, underpredict baryonic feedback as they cannot reproduce Sunyaev-Zel’dovich (SZ) measurements, which are more sensitive to the outskirts of haloes (Amodeo et al., 2021; Bigwood et al., 2024; Hadzhiyska et al., 2024). Although it might help resolve the $S_{8}$ tension, it raises a different problem because the X-ray and SZ measurements are no longer consistent with the same baryonic feedback model.
@@
Figure 1: Observational cosmological probes of the matter power spectrum and the wavenumber ($k)$, scale factor ($a$), and redshift ($z$) ranges to which they are approximately sensitive. From early to late Universe: cosmic microwave background (CMB, purple), CMB lensing ($\Phi\Phi$, light blue), the Lyman-alpha forest (Ly $\alpha$, orange), quasi-stellar objects (QSOs, red), galaxy clustering ($gg$, green), cosmic shear ($\gamma\gamma$, dark blue), and cluster abundance (CA, white). The black curve separates the linear and nonlinear collapse regimes. Constraining the matter power spectrum from different probes allows the localization of discrepancies in the two-dimensional ($k,a$) plane and provides insight into the physical origin of tensions between the measurements of different probes.
@@
In principle, we can distinguish tensions between measurements at different scales and times by directly constraining the matter power spectrum in the wavenumber ($k$) and scale factor ($a$) plane. Figure 1 illustrates the sensitivity of different observational cosmological probes in the ($k,a$) plane. The different probes are cosmic shear ($\gamma\gamma$), galaxy clustering ($gg$; e.g., Reid et al., 2010; Gil-Marín et al., 2016), CMB, CMB lensing ($\Phi\Phi$; e.g., Planck Collaboration et al., 2020b), cluster abundance (CA; e.g., Tegmark et al., 2004), quasi-stellar objects (QSOs; e.g., DESI Collaboration et al., 2024b), and the Lyman alpha forest (Ly $\alpha$; e.g., Palanque-Delabrouille et al., 2015; Karaçaylı et al., 2024). The squares correspond to probes where the emitted radiation is directly observed, whereas the ellipses correspond to processes happening along the line of sight between emission and observation. The black curve separates the linear and nonlinear collapse regimes. The estimate is obtained from class (Blas et al., 2011) by determining the wavenumber for which the nonlinear estimate differs by more than 5% from the linear prediction. The difference between the different cosmic times and scales probed by the CMB and cosmic shear, as described above, is clearly visible in the figure. Comparing the different probes and their cross-correlations allows the localization of the tension, which should inform us regarding the origin of the observed discrepancies.
@@
Preston et al. (2024) propose constraining the matter power spectrum from cosmic shear surveys. They estimate the constraining power of the matter power spectrum using synthetic cosmic shear measurements for the final data release of Stage III and Stage IV WL surveys. They find that while current cosmic shear surveys should only marginally be able to distinguish differences from a $\Lambda$CDM dark matter-only (DMO) model, next-generation surveys should provide clear constraints on the shape of the matter power spectrum at (mildly) nonlinear scales.
@@
Here, as a first step, we infer the best-fitting shape of the matter power spectrum using the latest cosmic shear measurement of Stage III WL surveys. We assume a simple parametric model that allows us to quantify the evolution of the matter power spectrum with $k$ and $a$ preferred by current data in an agnostic approach. The model allows the comparison of different surveys with different survey properties and does not assume a cosmological model for the matter power spectrum. We identify the scales to which the data is most sensitive, and test the impact of the assumed intrinsic alignment (IA) on the goodness of fit.
@@
In Sect. 2 we summarize the relevant WL theory to estimate the measured cosmic shear two-point statistics from the matter power spectrum. Section 3 describes the data, and the assumed parametric model is introduced in Sect. 4. In Sect. 5 we present the measurements of the best-fitting parametric model of the matter power spectrum. We also compare our results to the best-fitting $\Lambda$CDM predictions. The main findings are summarized in Sect. 6.
@@ 2 Theory
In this section we briefly summarize the theoretical framework for cosmic shear measurements as projected statistics of the matter power spectrum. Cosmic shear measures the systematic distortion of galaxy images from WL by the large-scale structure. The measured signal is comprised of the signal induced by weak gravitational lensing (G), the intrinsic alignment of the galaxies (I), and the cross-correlation of the two effects. The observed cosmic shear angular power spectrum ($\mathcal{C}^{(ij)}_{\epsilon\epsilon}$) between two populations of galaxies in different tomographic bins $i$ and $j$, as a function of multipole moment $\ell$, is given by
@@
Current cosmic shear surveys use $B$-mode statistics as null tests, and they find no evidence for nonzero $B$-modes or they limit their analysis to not include scales where the $B$-modes are nonzero due to unrecognized systematics. Here we assume that the $B$-modes are 0, and only use the $E$-mode signal such that $\mathcal{C}^{(ij)}_{\epsilon\epsilon}(\ell)=\mathcal{C}^{(ij)}_{\epsilon\epsilon,\mathrm{E}}(\ell)$. Assuming the modified Limber approximation, the angular power spectra are estimated from the matter power spectrum ($P_{\mathrm{m}}$) as (Limber, 1953; LoVerde & Afshordi, 2008)
@@
where $f_{\mathrm{K}}(\chi)$ is the angular diameter distance, $\chi$ the comoving distance, $z$ the redshift, and $\mathrm{a,b}\in{\mathrm{I},\mathrm{G}}$. The integral runs until the comoving line of sight horizon that is the edge of the galaxy survey ($\chi_{\mathrm{hor}}$). The WL kernel $W$ is given by (Kaiser, 1992)
@@
where $c$ is the speed of light, $a=1/(1+z)$ the scale factor, and $n_{\mathrm{s}}(\chi)$ the comoving distance source distribution that relates to the source redshift distribution as $n_{\mathrm{s}}(z)\mathrm{d}z=n_{\mathrm{s}}(\chi)\mathrm{d}\chi$. For the IA kernel, we assume the nonlinear alignment (NLA) model (Bridle & King, 2007)
@@
with $A_{\mathrm{IA}}$ the IA amplitude; $z_{\mathrm{piv}}$ a finite redshift pivot point; $\eta_{\mathrm{IA}}$ the redshift dependence exponent, which we set to 0; $C_{1}\rho_{\mathrm{cr}}\ =0.0134$; and $D$ the linear growth factor. We assume a flat Universe such that $f_{\mathrm{K}}(\chi)=\chi$. We adopt either $A_{\mathrm{IA}}=0$ or 1 to quantify whether including intrinsic alignment better fits the data.
We used three different cosmic shear observables in our analysis. Each observable is a projection of the angular power spectra, but is sensitive to different scales as they depend on different filter functions. The first observable is the shear two-point correlation functions (2PCFs, $\xi_{\mathrm{+/-}}$):
$\xi_{\mathrm{+/-}}^{(ij)}(\theta)=\int_{0}^{\infty}\frac{\mathrm{d}\ell\ell}{2\pi}\,\mathcal{J}_{0/4}(\ell\theta)\,\mathcal{C}^{(ij)}_{\epsilon\epsilon}(\ell).$ (5)
Here $\mathcal{J}_{0}$ and $\mathcal{J}_{4}$ are cylindrical Bessel functions of the first kind of the zeroth and fourth order, respectively. Second, we considered complete orthogonal sets of E/B-integrals (COSEBIs). An advantage of the COSEBIs is that nearly all cosmological information is contained in the first few $n$-modes, thus reducing the necessary computational time (Asgari et al., 2012). The COSEBIs ($E_{n}$) are estimated using (Schneider et al., 2010)
@@
where the weight function, $W_{n}(\ell)$, is a Hankel transform given by Eq. 9 in Asgari et al. (2021). Similar to Asgari et al. (2021), we used the first five COSEBIs $n$-modes.
Finally, we considered band power spectra ($\mathcal{C}_{\mathrm{E},l}$, band powers). This statistic is an angular average of the two-point correlation function that has a weak correlation between values at different wavenumbers and is given by (Schneider et al., 2002)
@@
where the normalisation $\mathcal{N}_{l}$ and kernels $W_{\mathrm{EE}}$ are respectively given by Eqs. 18 and 26 in Joachimi et al. (2021). We assume the best-fitting $\Lambda$CDM cosmology from Planck TT,TE,EE+lowE+lensing for the values of the cosmological parameters and to compute the redshift-comoving distance relation and linear growth factor (Planck Collaboration et al., 2020a).